A water tank has a hole 3.0 cm^2 in area located 2.0 m below the water surface. How many kilograms of water emerge from the hole per second?
1
Expert's answer
2021-08-11T09:54:04-0400

The velocity of water will be


"v=\sqrt{2gh}."

Find the volume:


"Q=vA=A\sqrt{2gh}."

Find the mass:


"m=Q\rho=\rho A\sqrt{2gh}=1000·3·10^{-4}\sqrt{2·2·9.8}=\\=1.88\text{ kg/s}."